# Last Iterate Convergence of Incremental Methods and Applications in Continual Learning

### v1

@@ Abstract
Incremental gradient methods and incremental proximal methods are a fundamental class of optimization algorithms used for solving finite sum problems, broadly studied in the literature. Yet, when it comes to their convergence guarantees, nonasymptotic (first-order or proximal) oracle complexity bounds have been obtained fairly recently, almost exclusively applying to the average iterate. Motivated by applications in continual learning, we obtain the first convergence guarantees for the last iterate of both incremental gradient and incremental proximal methods, in general convex smooth (for both) and convex Lipschitz (for the proximal variants) settings. Our oracle complexity bounds for the last iterate nearly match (i.e., match up to a square-root-log or a log factor) the best known oracle complexity bounds for the average iterate, for both classes of methods. We further obtain generalizations of our results to weighted averaging of the iterates with increasing weights, which can be seen as interpolating between the last iterate and the average iterate guarantees. Additionally, we discuss how our results can be generalized to variants of studied incremental methods with permuted ordering of updates. Our results generalize last iterate guarantees for incremental methods compared to state of the art, as such results were previously known only for overparameterized linear models, which correspond to convex quadratic problems with infinitely many solutions.
@@ 1 Introduction
Our main motivation for studying the last iterate convergence of incremental methods comes from applications in continual learning (CL). In particular, CL models a sequential learning setting, where a machine learning model gets updated over time, based on the changing or evolving distribution of the data passed to the learner. A major challenge in such dynamic learning settings is the degradation of model performance on previously seen data, known as the catastrophic forgetting (McCloskey and Cohen, 1989; Goodfellow et al., 2013), which has been well-documented in various empirical studies; see, e.g., recent surveys (De Lange et al., 2021; Parisi et al., 2019). On the theoretical front, however, much is still missing from the understanding of possibilities and limitations related to catastrophic forgetting, with results for basic learning settings being obtained very recently (Evron et al., 2022, 2023; Lin et al., 2023b; Goldfarb and Hand, 2023; Evron et al., 2024; Peng and Risteski, 2022; Peng et al., 2023; Chen et al., 2022; Cao et al., 2022; Balcan et al., 2015).
@@
where ${\bm{x}}_{n}\in{\mathbb{R}}^{d}$ is the model parameter vector used at time $n\in\mathbb{N}_{+}$. Such settings arise in applications that naturally undergo cyclic changes in the data/tasks, due to diurnal or seasonal cycles (e.g., in agriculture, forestry, e-commerce, astronomy, etc.).
Observe that (1.2) corresponds to the value of the objective function from (1.1) at the final iterate ${\bm{x}}_{KT}.$ Prior work (Evron et al., 2022) that obtained rigorous bounds for the forgetting (1.2) applied to the problems where each $f_{t}$ is a convex quadratic function minimized by the same ${\bm{x}}_{*}$ such that $f_{t}({\bm{x}}_{*})=f({\bm{x}}_{*})=0.$ By contrast, we consider more general convex functions that are either smooth or Lipschitz continuous, and make no assumption about ${\bm{x}}_{*}$ beyond being a minimizer of the (average) function $f$. Since we are not assuming that $f({\bm{x}}_{*})=0,$ our focus is on bounding the excess forgetting $f_{K}({\bm{x}}_{KT})-f({\bm{x}}_{*})$, which is equivalently the optimality gap for the last iterate in (1.1).
The method considered in Evron et al. (2022) minimized each component function exactly, outputting the solution closest to the previous iterate in each iteration, using implicit regularization properties of SGD. To obtain the results, it was then crucial that the component functions were quadratic (so that there is an explicit, closed-form solution for each subproblem) and that all component functions shared a nonempty set of minimizers with value zero (so that forgetting can be controlled despite aggressive adaption to the current task). Our work using a incremental proximal method instead considers explicit regularization to enforce closeness of models on differing tasks, which can potentially degrade the performance on the current task, but as a tradeoff can control forgetting and it addresses a much broader class of loss functions.
@@ 1.1 Contributions
Our main contributions can be summarized as follows.
@@
  • We provide the first oracle complexity guarantees for the last iterate of standard variants of the incremental gradient method with either deterministic or randomly permuted ordering of the updates, applied to convex $L$-smooth objectives. Up to a square-root-logarithm factor, our oracle complexity bounds – which are $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ for the deterministic variant and $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{\sqrt{TL}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ for the randomly permuted variant – match the best known oracle complexity bounds for these methods, previously known only for the (uniformly) average iterate (Mishchenko et al., 2020; Nguyen et al., 2021; Cai et al., 2023a). We further extend our results to increasing weighted averaging of iterates, which places more weight on more recent iterates, removing the excess square-root-log factor.
@@
  • We provide the first oracle complexity guarantees for the last iterate of the incremental proximal method, applied to convex and either smooth or Lipschitz-continuous objectives. When each component function is convex and $L$-smooth, we show that the incremental proximal method has the same $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ oracle complexity guarantee as the incremental gradient method. When component functions are convex and $G$-Lipschitz, our oracle complexity $\widetilde{\mathcal{O}}\Big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\Big{)}$ matches the best known oracle complexity bounds up to a logarithmic factor, which were previously known only for the (uniformly) average iterate (Bertsekas, 2011; Li et al., 2019). We further argue that our analysis can be extended to admit inexact proximal point evaluations.
  • We initiate the study of the proximal incremental method as a model of CL, corresponding to sequential ridge-regularized model training commonly used in practice. We discuss how the regularization parameter (or the parameter of proximal-point updates) affects the loss on the current learning task – determined by the regularization error – and (excess) catastrophic forgetting – corresponding to the optimality gap of the last iterate of the method.
@@ 1.2 Further Related Work
To remedy catastrophic forgetting, various empirical approaches have been proposed, corresponding to three main categories: (i) memory-based approaches, which store samples from previous tasks and reuse those data for training on the current task (Robins, 1995; Lopez-Paz and Ranzato, 2017; Rolnick et al., 2019); (ii) expansion-based approaches, which progressively expand the model size to accommodate new tasks (Rusu et al., 2016; Yoon et al., 2017); and (iii) regularization-based approaches, which regularize the loss of the current task to ensure the new model parameter vector is close to the old ones (Kirkpatrick et al., 2017); for a more complete survey, we refer to De Lange et al. (2021). Our focus in this work is on (i) and (ii).
On the theoretical front, the majority of existing work for the settings we consider has studied catastrophic forgetting for overparameterized linear models (Evron et al., 2022, 2023; Lin et al., 2023b; Goldfarb and Hand, 2023; Evron et al., 2024; Swartworth et al., 2024) or under the neural tangent kernel regime (Bennani et al., 2020; Doan et al., 2021). Another related line of work studied the task similarity in CL under the teacher-student learning framework (Asanuma et al., 2021; Lee et al., 2021). A recent work showed that an algorithm recursively minimizing the loss of the current task over the solution space of the previous task, named the ideal continual learner, never forgets, assuming all tasks share a common global minimizer (Peng et al., 2023).
@@
When it comes to the cyclic forgetting that this paper focuses on, theoretical results have only been established for linear models (Evron et al., 2022, 2023; Swartworth et al., 2024). In particular, the analysis for linear regression tasks in Evron et al. (2022); Swartworth et al. (2024) crucially relies on the exact minimization of each task using (S)GD to have closed-form updates between tasks, while Evron et al. (2023) uses alternating projections to analyze linear classification tasks. It is unclear how to extend either of these results to general convex loss functions that we consider.
@@
On a technical level, our results are most closely related to the literature on last iterate convergence guarantees for subgradient-based methods and stochastic gradient descent (SGD) and the literature on incremental gradient methods and shuffled SGD. For the former, we draw inspiration from the recent results (Zamani and Glineur, 2023; Liu and Zhou, 2023), which rely on a clever construction of reference points ${\bm{z}}_{k}$ with respect to which a gap quantity $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ gets bounded to deduce a bound for the optimality gap $f({\bm{x}}_{k})-f({\bm{x}}_{*})$ of the last iterate ${\bm{x}}_{*}$. For the latter line of work, we generalize the analysis used exclusively for the optimality gap of the (uniformly) average iterate (Mishchenko et al., 2020; Nguyen et al., 2021; Cha et al., 2023; Cai et al., 2023a) to control the gap-like quantities $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$, which require a more careful argument for controlling all error terms introduced by replacing ${\bm{x}}_{*}$ by ${\bm{z}}_{k-1}$ without introducing spurious unrealistic assumptions about the magnitudes of the component functions’ gradients. We finally note that both these related lines of work concern problems on which progress was made only in the very recent literature. In particular, while the oracle complexity upper bound for the average iterate of SGD in convex Lipschitz-continuous settings has been known for decades and its analysis is routinely taught in optimization and machine learning classes, there were no such results for the last iterate of SGD until 2013 (Shamir and Zhang, 2013) with improvements and generalizations to these results obtained as recently as in the past few months (Liu and Zhou, 2023). On the other hand, obtaining any nonasymptotic convergence guarantees for incremental gradient methods/shuffled SGD had remained open for decades (Bottou, 2009) until a recent line of work (Gürbüzbalaban et al., 2021; Shamir, 2016; Haochen and Sra, 2019; Nagaraj et al., 2019; Ahn et al., 2020; Rajput et al., 2020; Yun et al., 2022; Safran and Shamir, 2020). For smooth convex problems we consider, the convergence results were obtained only in the past few years (Mishchenko et al., 2020; Nguyen et al., 2021; Cha et al., 2023) and improved in Cai et al. (2023a) using a fine-grained analysis inspired by the recent advances in cyclic methods (Song and Diakonikolas, 2023; Cai et al., 2023b; Lin et al., 2023a). However, all those results are for the (uniformly) average iterate, while obtaining convergence results for the last iterate had remained open until this work.
@@ 1.3 Notation and Preliminaries
Throughout the paper, we consider the finite $d$-dimensional Euclidean space $(\mathbb{R}^{d},\|\cdot\|)$, where the norm $\|\cdot\|=\sqrt{\left\langle\cdot,\cdot\right\rangle}$ is induced by the standard inner product associated with the space. We let $\mathbb{N}_{+}$ be the set of positive natural numbers and denote $[T]:=\{1,2,\dots,T\}$. Given a proper, convex, lower semicontinuous function $f$, its proximal operator and Moreau envelope are defined by
$\mathrm{prox}_{\eta f}({\bm{x}})=\;$ $\operatorname*{arg\,min}_{{\bm{y}}\in{\mathbb{R}}^{d}}\Big{\{}\frac{1}{2\eta}\|{\bm{y}}-{\bm{x}}\|^{2}+f({\bm{y}})\Big{\}},$
$M_{\eta f}({\bm{x}})=\;$ $\min_{{\bm{y}}\in{\mathbb{R}}^{d}}\Big{\{}\frac{1}{2\eta}\|{\bm{x}}-{\bm{y}}\|^{2}+f({\bm{y}})\Big{\}},$
respectively, for a parameter $\eta>0$. The Moreau envelope can be seen as a smoothing of $f$, which is $\frac{1}{\eta}$-smooth with the gradient $\nabla M_{\eta f}({\bm{x}})=\frac{1}{\eta}({\bm{x}}-\mathrm{prox}_{\eta f}({\bm{x}}))\in\partial f(\mathrm{prox}_{\eta f}({\bm{x}}))$, where $\partial f(\cdot)$ denotes the subdifferential of $f$. To facilitate the discussion on inexact proximal point steps, following Lemaire (1992); Cominetti (1997); Alber et al. (1997); Salzo and Villa (2012); Villa et al. (2013), for a given error parameter $\epsilon\geq 0$ we further let $\partial_{\epsilon}f(\cdot)$ be the $\epsilon$-subdifferential of $f$ defined by (see the the textbook (Hiriart-Urruty and Lemarechal, 2013) for more details)
$\partial_{\epsilon}f({\bm{x}})=\{{\bm{g}}\in\mathbb{R}^{d}\mid f({\bm{y}})\geq f({\bm{x}})+\left\langle{\bm{g}},{\bm{y}}-{\bm{x}}\right\rangle-\epsilon,\forall{\bm{y}}\in\mathbb{R}^{d}\}.$
@@
Each component function $f_{t}$ is convex and there exists a minimizer ${\bm{x}}_{*}\in\operatorname*{arg\,min}_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}})$.
@@
Each $f_{t}$ is $G$-Lipschitz $(t\in[T])$, i.e., $|f_{t}({\bm{x}})-f_{t}({\bm{y}})|\leq G\|{\bm{x}}-{\bm{y}}\|$ for any ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$; thus $\|g_{t}({\bm{x}})\|\leq G$ for all $g_{t}({\bm{x}})\in\partial f_{t}({\bm{x}})$.
@@
Each $f_{t}$ is $L$-smooth $(t\in[T])$, i.e., $\|\nabla f_{t}({\bm{x}})-\nabla f_{t}({\bm{y}})\|\leq L\|{\bm{x}}-{\bm{y}}\|$ for any ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$.
We remark that 2 and 3 imply that $f$ is also $G$-Lipschitz and $L$-smooth, respectively. The assumption that all the component functions have the same Lipschitz or smoothness constant is possible to relax, in which case the resulting complexity bounds would scale with the average Lipschitz/smoothness constant. This can be done using the techniques from our prior work (Cai et al., 2023a). However, to avoid complicating the analysis and to keep the focus on the intricacies of the last iterate convergence, here we maintain the simpler setup under 2 and 3, and defer generalizations to the setting of distinct component Lipschitz/smoothness constants to later extended version of this paper.
When $f$ is $L$-smooth and convex, we will often make use of the following standard inequality that fully characterizes the class of $L$-smooth convex functions:
@@
Finally, when each $f_{t}$ is smooth, we assume bounded variance at ${\bm{x}}_{*}$, same as prior work (Mishchenko et al., 2020; Nguyen et al., 2021; Tran et al., 2021, 2022; Cai et al., 2023a).
@@
2 Incremental Gradient Descent and Shuffled SGD
In this section, we analyze the oracle complexity for the last iterate of incremental gradient descent (IGD), assuming component functions are smooth and convex. In the context of CL, this corresponds to a simplified setup where the learner incrementally performs a single gradient step on each task and cyclically replays the $T$ tasks. Nevertheless, this setup serves as a warmup to the proximal setup we discuss in the next section. Additionally, it is of independent interest as incremental gradient methods are widely used in the optimization and machine learning literature, where despite the lack of prior theoretical justification, it is typically the last iterate that gets output by the algorithm in practice.
We summarize the IGD method in Algorithm 1, assuming the incremental order $1,2,\dots,T$ in each epoch for simplicity and without loss of generality. The oracle complexity for the (uniformly) average iterate of IGD has been shown to be ${\mathcal{O}}(\frac{TL}{\epsilon}+\frac{T\sqrt{L}\sigma_{*}}{\epsilon^{3/2}})$ for an $\epsilon$-optimality gap (Mishchenko et al., 2020; Cai et al., 2023a) under the same assumptions we make here (1, 3, and 4), while, as discussed before, there were no guarantees for either the last iterate or even a weighted average of the iterates. The main result of this section is that the same oracle complexity applies to the last iterate of IGD, up to a square-root-log factor. We then further generalize this result to weighted averages of iterates with increasing weights and to variants with the randomly permuted order of cyclic updates.

### v2

@@ Abstract
Incremental gradient and incremental proximal methods are a fundamental class of optimization algorithms used for solving finite sum problems, broadly studied in the literature. Yet, without strong convexity, their convergence guarantees have primarily been established for the ergodic (average) iterate. Motivated by applications in continual learning, we obtain the first convergence guarantees for the last iterate of both incremental gradient and incremental proximal methods, in general convex smooth (for both) and convex Lipschitz (for the proximal variants) settings. Our oracle complexity bounds for the last iterate nearly match (i.e., match up to a square-root-log or a log factor) the best known oracle complexity bounds for the average iterate, for both classes of methods. We further obtain generalizations of our results to weighted averaging of the iterates with increasing weights and for randomly permuted ordering of updates. We study incremental proximal methods as a model of continual learning with generalization and argue that large amount of regularization is crucial to preventing catastrophic forgetting. Our results generalize last iterate guarantees for incremental methods compared to state of the art, as such results were previously known only for overparameterized linear models, which correspond to convex quadratic problems with infinitely many solutions.
@@ 1 Introduction
Our main motivation for studying the last iterate convergence of incremental methods comes from applications in continual learning (CL). In particular, CL models a sequential learning setting, where a machine learning model gets updated over time, based on the changing or evolving distribution of the data passed to the learner. A major challenge in such dynamic learning settings is the degradation of model performance on previously seen data, known as the catastrophic forgetting (McCloskey and Cohen, 1989; Goodfellow et al., 2013), which has been well-documented in various empirical studies; see, e.g., recent surveys (De Lange et al., 2021; Parisi et al., 2019). On the theoretical front, however, much is still missing from the understanding of possibilities and limitations related to catastrophic forgetting, with results for basic learning settings being obtained only very recently (Evron et al., 2022, 2023; Lin et al., 2023b; Goldfarb and Hand, 2023; Goldfarb et al., 2024; Peng and Risteski, 2022; Peng et al., 2023; Chen et al., 2022; Cao et al., 2022; Balcan et al., 2015).
@@
where ${\bm{x}}_{n}\in{\mathbb{R}}^{d}$ is the model parameter vector used at time $n\in\mathbb{N}_{+}$. Such settings arise in applications that naturally undergo cyclic changes in the data/tasks, due to diurnal or seasonal cycles (e.g., in agriculture, forestry, e-commerce, astronomy, etc.). Forgetting is catastrophic if $f_{K}({\bm{x}}_{KT})\overset{K\rightarrow\infty}{\nrightarrow}0$.
Observe that Eq. (1.2) corresponds to the value of the objective function from Eq. (1.1) at the final iterate ${\bm{x}}_{KT}.$ Prior work (Evron et al., 2022) that obtained rigorous bounds for the forgetting Eq. (1.2) applied to the problems where each $f_{t}$ is a convex quadratic function minimized by the same ${\bm{x}}_{*}$ such that $f_{t}({\bm{x}}_{*})=f({\bm{x}}_{*})=0.$ By contrast, we consider more general convex functions that are either smooth or Lipschitz continuous, and make no assumption about ${\bm{x}}_{*}$ beyond being a minimizer of the (average) function $f$. Since we are not assuming that $f({\bm{x}}_{*})=0,$ our focus is on bounding the excess forgetting $f_{K}({\bm{x}}_{KT})-f({\bm{x}}_{*})$, which is equivalently the optimality gap for the last iterate in Eq. (1.1).
The method considered in Evron et al. (2022) minimized each component function exactly, outputting the solution closest to the previous iterate in each iteration, using implicit regularization properties of SGD. To obtain the results, it was then crucial that the component functions were quadratic (so that there is an explicit, closed-form solution for each subproblem) and that all component functions shared a nonempty set of minimizers with value zero (so that forgetting can be controlled despite aggressive adaption to the current task). Our work using the incremental proximal method instead considers explicit regularization to enforce closeness of models on differing tasks, which can potentially degrade the performance on the current task, but as a tradeoff can control forgetting and it addresses a much broader class of loss functions.
@@ 1.1 Contributions
Our main contributions can be summarized as follows, where $\sigma_{*}^{2}:=\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{*})\|^{2}$ denotes the gradient variance at the optimum. The quantity $\sigma_{*}^{2}$ is intrinsic to oracle complexity of incremental methods (Mishchenko et al., 2020; Nguyen et al., 2021; Cai et al., 2023a; Cha et al., 2023).
Last iterate convergence of Incremental Gradient Descent (IGD).
@@
We provide the first oracle complexity guarantees for the last iterate of standard variants of IGD with either deterministic or randomly permuted ordering of the updates, applied to convex $L$-smooth objectives. Up to a square-root-log factor, our oracle complexity bounds in Theorem 2.3 and Corollary 2.5 – which are $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ for the deterministic variant and $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{\sqrt{TL}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ for the randomly permuted variant – match the best known oracle complexity bounds for these methods, previously known only for the (uniformly) average iterate (Mishchenko et al., 2020; Nguyen et al., 2021; Cai et al., 2023a; Cha et al., 2023). We further extend our results to increasing weighted averaging of the iterates in Corollary 2.4, which places more weight on the more recent iterates, removing the excess square-root-log factor in the resulting oracle complexity bound.
Last iterate convergence of Incremental Proximal Method (IPM).
@@
We provide the first oracle complexity guarantees for the last iterate of IPM applied to convex and either smooth or Lipschitz-continuous objectives. When each component function is convex and $L$-smooth, we show (in Theorem 3.1) that IPM has the same $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ oracle complexity guarantee as IGD. This result is new for any variant of this method – with average or last iterate as its output. When component functions are convex and $G$-Lipschitz, our oracle complexity $\widetilde{\mathcal{O}}\Big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\Big{)}$ in Theorem 3.3 matches the best known oracle complexity bound up to a log factor, which was previously known only for the (uniformly) average iterate (Bertsekas, 2011; Li et al., 2019). We further argue (in Corollary 3.4 and Corollary 3.5) that for both settings our analysis can be extended to admit inexact proximal point evaluations – an important setting not addressed by prior work on general IPM.
IPM as a model of CL.
We initiate the study of IPM as a model of CL, corresponding to sequential ridge-regularized model training commonly used in practice. On the positive side, our last-iterate convergence results for IPM in Theorem 3.1 and Theorem 3.3 demonstrate that forgetting (corresponding to the optimality gap at the last iterate) can be effectively controlled if the amount of employed regularization is sufficiently high. On the negative side, we show that for any constant amount of regularization, forgetting is always catastrophic, even for least squares problems. In particular, we provide a univariate quadratic example such that for any constant regularization parameter, the asymptotic limit of (excess) forgetting is non-zero. Further, we demonstrate that for forgetting to be made smaller than some target $\epsilon,$ the regularization must be sufficiently high and depend polynomially on $1/\epsilon,\,T,$ and $\sigma_{*}.$ These results are summarized in Theorem 3.2 and highlight the limitations of regularization as a black-box tool for controlling forgetting in CL.
@@ 1.2 Further related work
To remedy catastrophic forgetting, various empirical approaches have been proposed, and this work is most closely related to (i) memory-based approaches, which store samples from previous tasks and reuse those data for training on the current task (Robins, 1995; Lopez-Paz and Ranzato, 2017; Rolnick et al., 2019) and (ii) regularization-based approaches, which regularize the loss of the current task to ensure the new model parameter vector is close to the prior ones (Kirkpatrick et al., 2017); for a more complete survey, we refer to De Lange et al. (2021). On the theoretical front, the results on the forgetting for cyclic replaying of tasks considered in our work have only been established for linear models (Evron et al., 2022, 2023; Swartworth et al., 2024). In particular, the analysis for linear regression tasks in Evron et al. (2022); Swartworth et al. (2024) crucially relies on the exact minimization of each task using (S)GD to have closed-form updates between tasks, while Evron et al. (2023) uses alternating projections to analyze linear classification tasks. It is unclear how to extend either of these results to general convex loss functions that we consider.
@@
On a technical level, our results are most closely related to 1) the literature on last iterate convergence guarantees for subgradient-based methods and stochastic gradient descent (SGD) and 2) the literature on incremental gradient methods and shuffled SGD. For 1), we draw inspiration from the recent results (Zamani and Glineur, 2023; Liu and Zhou, 2023), which rely on a clever construction of reference points ${\bm{z}}_{k}$ with respect to which a gap quantity $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ gets bounded to deduce a bound for the optimality gap $f({\bm{x}}_{k})-f({\bm{x}}_{*})$ of the last iterate ${\bm{x}}_{*}$. For the latter line of work 2), we generalize the analysis used exclusively for the optimality gap of the (uniformly) average iterate (Mishchenko et al., 2020; Nguyen et al., 2021; Cha et al., 2023; Cai et al., 2023a) to control the gap-like quantities $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$, which require a more careful argument for controlling all error terms introduced by replacing ${\bm{x}}_{*}$ by ${\bm{z}}_{k-1}$ without introducing spurious unrealistic assumptions about the magnitudes of the component functions’ gradients. We finally note that both these related lines of work concern problems on which progress was made only in the very recent literature. In particular, while the oracle complexity upper bound for the average iterate of SGD in convex Lipschitz-continuous settings has been known for decades and its analysis is routinely taught in optimization and machine learning classes, there were no such results for the last iterate of SGD until 2013 (Shamir and Zhang, 2013) with improvements and generalizations to these results obtained as recently as in the past year (Liu and Zhou, 2023). Regarding 2), obtaining any nonasymptotic convergence guarantees for incremental gradient methods/shuffled SGD had remained open for decades (Bottou, 2009) until a recent line of work (Gürbüzbalaban et al., 2021; Shamir, 2016; Haochen and Sra, 2019; Nagaraj et al., 2019; Ahn et al., 2020; Rajput et al., 2020; Yun et al., 2022; Safran and Shamir, 2020). For nonconvex problems, Yu and Li (2023) proved the high-probability last iterate guarantee for shuffled SGD with stopping criteria, which is technically disjoint from our work. For smooth convex problems we consider, the convergence results were obtained only in the past few years (Mishchenko et al., 2020; Nguyen et al., 2021; Cha et al., 2023) and improved in Cai et al. (2023a) using a fine-grained analysis inspired by the recent advances in cyclic methods (Song and Diakonikolas, 2023; Cai et al., 2023b; Lin et al., 2023a). However, all those results are for the (uniformly) average iterate, while obtaining convergence results for the last iterate had remained open.
Concurrent independent work.
An independent and concurrent work to ours (Liu and Zhou, 2024) studied the last-iterate convergence of shuffled SGD for composite (strongly) convex smooth/Lipschitz optimization. For the same problems as studied in our Section 2, they obtained the same convergence results. The remaining results in Liu and Zhou (2024) and our work are not directly comparable, as the motivation for the two works and the studied settings are different. In particular, the focus of Liu and Zhou (2024) is on the last iterate convergence of shuffled SGD, and thus they study it in depth, considering different Lipschitz/smoothness constants for component functions, strong convexity, and composite settings. Our focus on the other hand is on settings relevant to continual learning, and thus we additionally consider the convergence of a weighted average of iterates and put more weight on the incremental proximal method, which were not considered in Liu and Zhou (2024). It is of note that while Liu and Zhou (2024) used proximal steps to handle the nonsmooth portion of the objective in their composite setting, the proximal maps are not applied component-wise, but at the end of a cycle, to only one (regularizer) function (e.g., to handle constraints or joint regularization).
@@ 1.3 Notation and preliminaries
[…]
$\mathrm{prox}_{\eta f}({\bm{x}})=\operatorname*{arg\,min}_{{\bm{y}}\in{\mathbb{R}}^{d}}\Big{\{}\frac{1}{2\eta}\|{\bm{y}}-{\bm{x}}\|^{2}+f({\bm{y}})\Big{\}},\;M_{\eta f}({\bm{x}})=\min_{{\bm{y}}\in{\mathbb{R}}^{d}}\Big{\{}\frac{1}{2\eta}\|{\bm{y}}-{\bm{x}}\|^{2}+f({\bm{y}})\Big{\}},$
respectively, for a parameter $\eta>0$. The Moreau envelope is $\frac{1}{\eta}$-smooth with the gradient $\nabla M_{\eta f}({\bm{x}})=\frac{1}{\eta}({\bm{x}}-\mathrm{prox}_{\eta f}({\bm{x}}))\in\partial f(\mathrm{prox}_{\eta f}({\bm{x}}))$, where $\partial f(\cdot)$ denotes the subdifferential of $f$.
@@
Each $f_{t}$ is convex and there exists a minimizer ${\bm{x}}_{*}\in\operatorname*{arg\,min}_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}})$.
@@
Each $f_{t}$ is $G$-Lipschitz, i.e., $|f_{t}({\bm{x}})-f_{t}({\bm{y}})|\leq G\|{\bm{x}}-{\bm{y}}\|$ for any ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$; thus $\|g_{t}({\bm{x}})\|\leq G$ for all $g_{t}({\bm{x}})\in\partial f_{t}({\bm{x}})$.
@@
Each $f_{t}$ is $L$-smooth, i.e., $\|\nabla f_{t}({\bm{x}})-\nabla f_{t}({\bm{y}})\|\leq L\|{\bm{x}}-{\bm{y}}\|$ for any ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$.
We remark that Assumptions 2 and 3 imply that $f$ is also $G$-Lipschitz and $L$-smooth, respectively, These two assumptions can also be generalized to be with distinct Lipschitz/smoothness constants, and our results would scale with the average Lipschitz/smoothness constant using the techniques from Cai et al. (2023a), which we omit to keep the focus on the intricacies of the last iterate convergence. When $f$ is $L$-smooth and convex, we will often make use of the following standard inequality that fully characterizes the class of $L$-smooth convex functions:
@@
Finally, when each $f_{t}$ is smooth, we assume bounded variance at ${\bm{x}}_{*}$, same as all prior work that considered the same settings of IGD/shuffled SGD as we do (Mishchenko et al., 2020; Nguyen et al., 2021; Tran et al., 2021, 2022; Cai et al., 2023a).
@@
2 Last Iterate Convergence of Incremental Gradient Descent
In this section, we introduce our techniques for analyzing the last iterate guarantee and bound the oracle complexity for the last iterate of incremental gradient descent (IGD), assuming component functions are smooth and convex. In the context of CL, this corresponds to a simplified setup where the learner incrementally performs a single gradient step on each task and cyclically replays the $T$ tasks. Nevertheless, this setup serves as a warmup to the proximal setup we discuss in the next section. Additionally, it is of independent interest as incremental gradient methods are widely used in the optimization and machine learning literature, where despite the lack of prior theoretical justification, it is typically the last iterate that gets output by the algorithm in practice.
We summarize the IGD method in Alg. 1, assuming the incremental order $1,2,\dots,T$ in each epoch for simplicity and without loss of generality. The oracle complexity for the (uniformly) average iterate of IGD has been shown to be ${\mathcal{O}}(\frac{TL}{\epsilon}+\frac{T\sqrt{L}\sigma_{*}}{\epsilon^{3/2}})$ for an $\epsilon$-optimality gap (Mishchenko et al., 2020; Cai et al., 2023a) under the same assumptions we make here (Assumptions 13, and 4), while, as discussed before, there were no guarantees for either the last iterate or even a weighted average of the iterates. The main result of this section is that the same oracle complexity applies to the last iterate of IGD, up to a square-root-log factor. We then further generalize this result to weighted averages of iterates with increasing weights and to variants with randomly permuted order of cyclic updates.